Determine the price of the security:

A share of stock that pays a current dividend of $2.94. The return on equity is 11.2% and total earnings for the past year is $6.29. Shareholders required rate of return is 12.6%.
Answer #1

Dividend payout ratio = DPS/EPS = 2.94/6.29

=0.4674

Growth rate=ROE*(1-payout ratio)
growth rate=11.2*(1-0.4674)
growth rate = 5.97
Price = recent dividend* (1 + growth rate )/(cost of equity - growth rate)
Price = 2.94 * (1+0.0597) / (0.126 - 0.0597)
Price = 46.99